# The infrared luminosity of retired and post-starburst galaxies: A cautionary tale for star formation rate measurements

### v1

@@ 6 Summary
  • We use toy models to show that the excess $L_{\mathrm{TIR}}$ is predominantly due to the increasing dominance of old, low mass stars heating the dust, and not due to recent star formation history. However, the rare population of galaxies with ongoing, yet rapidly declining, star formation rates (young post-starburst galaxies) may have slightly enhanced $L_{\mathrm{TIR}}$ simply due to the longer averaging time of $L_{\mathrm{TIR}}$ than $L_{\mathrm{H}\,\alpha}$ when converted into an SFR.
  • Recently and rapidly quenched post-starburst galaxies do not show any evidence of hosting dust-obscured starbursts. Their high $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratios are entirely consistent with being caused by the above processes.
An improved understanding of heating and ionisation mechanisms in retired and post-starburst galaxies could be gained from spatially resolved observations, both in the mid- and far-IR and with optical integral field spectroscopy. Sensitive radio observations would also provide additional assurance that significant amounts of dust obscured star formation was not being missed from the optical census of the Universe. However, our results indicate that the high fraction of rapidly quenched galaxies detected in optical surveys is not a result of high levels of contamination from dust-obscured starbursts, and therefore understanding rapid quenching pathways to quiescence remains of strong importance to understanding galaxy evolution.
@@ Acknowledgements
We would like to thank Rob Kennicutt and Desika Narayanan for their help and advice in interpreting the results, and Kenneth Duncan, Ho-Seong Hwang and Leah Morabito for sharing data and information on datasets during preliminary investigations carried out prior to this paper being written.
@@ Appendix A Star formation histories for toy SED models
Figure 4 shows the average starlight fitted star formation histories of our SF, post-starburst and retired galaxy subsamples divided into five stellar mass bins (Cid Fernandes et al., 2005). These average star formation histories are used to calculate $L_{\mathrm{TIR}}$ and $L_{\mathrm{H}\,\alpha}$ in our toy models, as described in Section 5. We excluded the two lowest mass bins for the retired and post-starburst samples, due to the low number of galaxies or very low predicted luminosities. The visible bump at 1 Gyr lookback time in the star-forming galaxy star formation histories, and to a lesser extent in the retired galaxy star formation histories, is caused by imperfections in the stellar population models. The youngest age bin is similarly affected and is removed from the analysis, as described in Section 5 (Ocvirk, 2010; Cid Fernandes et al., 2011). The logarithmic time axis is scaled deliberately to highlight the recent past history of these galaxies, which is the important part for our analysis. The post-starburst galaxies have had a notably stronger burst of star formation between 500 and 1.5 Gyr in lookback time than caused by the model imperfections, but have experienced little star formation since then. The quiescent galaxies have been quiescent for at least 1 Gyr.
These star formation histories show that the star formation rate of all of these populations has been very stable for the last 100 Myr. This is important for our discussion of the cause of the excess $L_{\mathrm{TIR}}$ in the retired and post-starburst galaxies in Section 5.
Figure 4: Average star formation histories for the retired, post-starburst (PSB) and star-forming (SF) subsamples in stellar mass bins, used in the toy models in Section 5. The amount of mass formed as a function of stellar population age $t_{\star}$ is represented by the initial mass-fraction population vector $\mu_{\mathrm{ini}}$.

### v2

@@ 6 Summary
  • We use toy models to show that the excess $L_{\mathrm{TIR}}$ is predominantly due to the increasing dominance of old, low mass stars heating the dust, and not due to a recent decline in star formation rate, combined with the assumption of constant star formation rates in the calibration conversions of luminosity into SFR. However, the rare population of galaxies with ongoing, yet rapidly declining, star formation rates (young post-starburst galaxies) may have slightly enhanced $L_{\mathrm{TIR}}$ simply due to the longer averaging time of $L_{\mathrm{TIR}}$ than $L_{\mathrm{H}\,\alpha}$ when converted into an SFR.
  • Recently and rapidly quenched local post-starburst galaxies do not show any evidence of hosting dust-obscured starbursts. Their high $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratios are entirely consistent with being caused by the above processes.
An improved understanding of heating, ionisation and dust destruction mechanisms in retired and post-starburst galaxies could be gained from spatially resolved observations, both in the mid- and far-IR and with optical integral field spectroscopy. Sensitive radio observations would also provide additional assurance that significant amounts of dust obscured star formation was not being missed from the optical census of the Universe. However, our results indicate that the high fraction of rapidly quenched galaxies detected in optical surveys is not a result of high levels of contamination from dust-obscured starbursts, and therefore understanding rapid quenching pathways to quiescence remains of strong importance to understanding galaxy evolution.
@@ Acknowledgements
We would like to thank Rob Kennicutt and Desika Narayanan for their help and advice in interpreting the results, and Kenneth Duncan, Ho-Seong Hwang and Leah Morabito for sharing data and information on datasets during preliminary investigations carried out prior to this paper being written. We would also like to thank Ian Smail and the anonymous OJAp referees who provided thoughtful and helpful feedback that improved this manuscript.
@@ Appendix A Star formation histories for toy SED models
Figure 5: Average star formation histories for the retired, post-starburst (PSB) and star-forming (SF) subsamples in stellar mass bins, used in the toy models in Section 5. The cumulative star formation rate is shown as function of the stellar population age $t_{\star}$.
@@
Figure 5 shows the average starlight fitted star formation histories (SFH) of our star-forming, post-starburst and retired galaxy subsamples split into five stellar mass bins. These are displayed in cumulative form, in order to best highlight the differences between the different subsamples. These average star formation histories are used to calculate $L_{\mathrm{TIR}}$ and $L_{\mathrm{H}\,\alpha}$ in our toy models, as described in Section 5. We excluded the two lowest mass bins for the retired and post-starburst samples, due to the low number of galaxies or very low predicted luminosities.
We can see that retired galaxies have been quiescent for at least 1 Gyr, at all stellar masses. Post-starburst galaxies experienced a significant burst of star formation between 500 and 1.5 Gyr ago and have undergone minimal star formation since. Low mass star-forming galaxies show substantial star formation within the past 1 Gyr, while higher mass galaxies formed an increasingly insignificant proportion of their mass recently. These star formation histories demonstrate that the star formation rate of all galaxies with stellar masses larger than $10^{9}$, as in our samples (Figure 1), has been very stable for the last 100 Myr. This is important for our discussion of the cause of the excess $L_{\mathrm{TIR}}$ in the retired and post-starburst galaxies in Section 5.
Our conversion from star formation history into $L_{\mathrm{TIR}}$/ $L_{\mathrm{H}\,\alpha}$ depends on many properties of the stellar population models and dust attenuation curve, and the magnitude of the effect in Figure 4 should only be taken as indicative. We investigated the impact of changing the assumed attenuation curve to a shallower Charlot & Fall (2000) dust law, with $q_{\lambda}=(\lambda/{5500\,\mathrm{\AA}})^{-0.7}$. This results in a modest increase in the predicted $\log L_{\mathrm{TIR}}$ of 0.16 dex for the star-forming and of 0.08 dex for the post-starburst and retired galaxies. More recent stellar population models than the BC03 models used here yield slightly different SFHs when fit to optical spectra, e.g. a newer version of the Bruzual & Charlot models which include the new MILES stellar spectral library result in smoother SFHs (e.g. Pawlik et al., 2018). Improvements in UV modelling post-Bruzual & Charlot (2003) may slightly shift our results, but the key challenge lies in accurately characterizing young stellar populations. From a technical perspective, incorporating UV constraints into optical spectral synthesis can help to better constrain these populations (e.g. López Fernández et al., 2018; Werle et al., 2019). However, even the latest stellar libraries face limitations, particularly in their coverage of the crucially important hot young stars. Excluding stellar populations younger than $10^{8}$ years from our calculations is not ideal, as these do contribute some light to the galaxy SED, if not much mass. However, excluding these ages is a common approach taken to account for these limitations, with some stellar population models only starting at 30–60 Myr (e.g. Vazdekis et al., 2010).